# HAWX: A Hardware-Aware FrameWork for Fast and Scalable ApproXimation of DNNs

### v1

@@ I Introduction
Moreover, few approaches explicitly account for the diversity of DNN hardware accelerators. Spatial architectures like dataflow accelerators rely on massive parallelism and data reuse [3], while temporal architectures, such as FPGA-based soft processors (e.g., FGPU), depend on flexible time-multiplexed execution [2]. Efficient AxC integration must consider these architectural differences, yet most existing frameworks are hardware-agnostic or target only a narrow class of architectures.
@@ II-C Hardware-Aware Approximate Configurations
After the boundary hawxhawx scores are determined, hardware information is incorporated to map approximate blocks to the target platform. The hardware platform is first identified (Algorithm 1, Line 1), such as a spatial (e.g., dataflow) or temporal accelerator (e.g., FGPU). For each platform, architectural parameters are provided to guide block assignment and estimation of hardware cost.
@@ III-A Experimental Setup
Software-based evaluations are carried out on a workstation equipped with an Intel Core i7-7500U CPU, 16 GB of RAM, and an NVIDIA GeForce 940MX GPU. To further validate hardware-level behavior, selected approximate configurations for both spatial and temporal DNN accelerators are deployed on an FPGA platform (Xilinx Zynq UltraScale+). Hardware synthesis and implementation are performed using Xilinx Vivado and Vitis HLS 2020.2. As an example of spatial architectures, dataflow-style accelerators are generated by first converting the approximated DNN code into synthesizable C using DeepAxe [14], followed by hardware synthesis. As an example of temporal architectures, FGPU-style soft processors are employed, in which computations are mapped to several compute vectors on the FPGA fabric [2]. Resource allocation and scheduling for both hardware contexts are managed through the HAWX framework.
@@ III-B Results and Discussion
Figure 5 further validates these findings, presenting accuracy–power–area trade-offs for LeNet-5 (blue) and VGG-11 (red) on both spatial (dataflow) and temporal (FGPU) architectures. The synthesized results include both layer-wise and filter-wise configurations for both architectures. For the dataflow accelerator, layers are mapped directly onto the hardware, exposing layer-level parallelism (circles). The dataflow configurations correspond to those reported in Table II, with the star marker denoting the exact baseline configuration.
In contrast, temporal execution on FGPU accelerators (squares) leverages CVs, each comprising multiple PEs with distinct Local IDs (LIDs). Approximation can be selectively applied across LIDs within a CV or across CVs, enabling heterogeneous execution. In these experiments, FGPU configurations employ L2N and KV8 multipliers, with one CV instantiated for LeNet-5 and eight CVs for VGG-11. During execution, the layer or filter is partitioned into work-items; these are grouped into work-groups, which the Work-Group Dispatcher schedules onto available CVs. Each work item is mapped to a specific LID according to the chosen configuration. The results confirm that HAWX-selected AxC configurations achieve substantial reductions in power and area while maintaining accuracy, highlighting gains across both hardware types.
90909292949496960.970.970.980.980.990.99110.90.911Accuracy (%)PowerAreaLeNet-5 (FGPU)VGG-11 (FGPU)LeNet-5 (Dataflow)VGG-11 (Dataflow)
Figure 5: Trade-off between accuracy, power, and area for dataflow and FGPU architectures.

### v2

@@ I Introduction
Moreover, few approaches explicitly account for the diversity of DNN hardware accelerators. Spatial architectures like dataflow accelerators rely on massive parallelism and data reuse [3], while temporal architectures, such as FPGA-based soft processors (e.g., PERUN), depend on flexible time-multiplexed execution [2]. Efficient AxC integration must consider these architectural differences, yet most existing frameworks are hardware-agnostic or target only a narrow class of architectures.
@@ II-C Hardware-Aware Approximate Configurations
After the boundary hawxhawx scores are determined, hardware information is incorporated to map approximate blocks to the target platform. The hardware platform is first identified (Algorithm 1, Line 1), such as a spatial (e.g., dataflow) or temporal accelerator (e.g., PERUN). For each platform, architectural parameters are provided to guide block assignment and estimation of hardware cost.
@@ III-A Experimental Setup
Software-based evaluations are carried out on a workstation equipped with an Intel Core i7-7500U CPU, 16 GB of RAM, and an NVIDIA GeForce 940MX GPU. To further validate hardware-level behavior, selected approximate configurations for both spatial and temporal DNN accelerators are deployed on an FPGA platform (Xilinx Zynq UltraScale+). Hardware synthesis and implementation are performed using Xilinx Vivado and Vitis HLS 2020.2. As an example of spatial architectures, dataflow-style accelerators are generated by first converting the approximated DNN code into synthesizable C using DeepAxe [14], followed by hardware synthesis. As an example of temporal architectures, PERUN-style soft processors are employed, in which computations are mapped to several compute vectors on the FPGA fabric. PERUN is a low-power run-time adaptive multi-threaded scalable neural network accelerator designed by the NEUROFABRIX company. Resource allocation and scheduling for both hardware contexts are managed through the HAWX framework.
@@ III-B Results and Discussion
Figure 5 further validates these findings, presenting accuracy–power–area trade-offs for LeNet-5 (blue) and VGG-11 (red) on both spatial (dataflow) and temporal (PERUN) architectures. The synthesized results include both layer-wise and filter-wise configurations for both architectures. For the dataflow accelerator, layers are mapped directly onto the hardware, exposing layer-level parallelism (circles). The dataflow configurations correspond to those reported in Table II, with the star marker denoting the exact baseline configuration.
In contrast, temporal execution on the PERUN accelerator (squares) leverages CVs, each comprising multiple PEs with distinct Local IDs (LIDs). Approximation can be selectively applied across LIDs within a CV or across CVs, enabling heterogeneous execution. In these experiments, PERUN configurations employ L2N and KV8 multipliers, with one CV instantiated for LeNet-5 and eight CVs for VGG-11. During execution, the layer or filter is partitioned into work-items; these are grouped into work-groups, which the Work-Group Dispatcher schedules onto available CVs. Each work item is mapped to a specific LID according to the chosen configuration. The results confirm that HAWX-selected AxC configurations achieve substantial reductions in power and area while maintaining accuracy, highlighting gains across both hardware types.
90909292949496960.970.970.980.980.990.99110.90.911Accuracy (%)PowerAreaLeNet-5 (PERUN)VGG-11 (PERUN)LeNet-5 (Dataflow)VGG-11 (Dataflow)
Figure 5: Trade-off between accuracy, power, and area for dataflow and PERUN architectures.